OpenStudy (anonymous):

Begin with the equation 4= 2^2. Raise each side to the exponent 1/2 and use the power of a power law to simplify the expression on the right side.

OpenStudy (anonymous):

\[4=2^2\] \[4^{\frac{1}{2}}=(2^2)^{\frac{1}{2}}\] \[4^{\frac{1}{2}}=2^{2\times \frac{1}{2}}\] \[4^{\frac{1}{2}}=2^1=2\]

OpenStudy (anonymous):

is this supposed to convince you that \[4^{\frac{1}{2}}=\sqrt{4}\]?

OpenStudy (anonymous):

yes, it is.